# Measuring Uncertainty Calibration

### v2

@@ Abstract
We make two contributions to the problem of estimating the $L_{1}$ calibration error of a binary classifier from a finite dataset. First, we provide an upper bound for any classifier where the calibration function has bounded variation. Second, we provide a method of modifying any classifier so that its calibration error can be upper bounded efficiently without significantly impacting classifier performance and without any restrictive assumptions. All our results are non-asymptotic and distribution-free. We conclude by providing advice on how to measure calibration error in practice. Our methods yield practical procedures that can be run on real-world datasets with modest overhead.
@@ 3 Structure of Results
Here, we used the symbol $B$ to denote the concentration term in the Bernstein bound. We could have equally used Hoeffding inequality here, but Bernstein is slightly sharper.
@@ 4 Measuring Calibration Under Bounded Variation
To formalize our assumption on $\eta$, we will need the concept of a Bounded Variation space (DeVore & Lorentz, 1993). For a function $f$, we can define total variation on the interval II as
@@
where the supremum is taken over the set of all finite sequences $s_{1},s_{2},\dots,s_{n}$ contained in the interval II. We say that the function $f$ is of bounded variation, written $f\in\text{BV}([0,1])$ when $\text{TV}(f,[0,1])<\infty$. Our theoretical work in this section is based on the assumption that the calibration function $\eta$ satisfies $\text{TV}(\eta,[0,1])\leq V$ for some known constant $V$.
@@
Under the bounded variation assumption, we want to reconstruct $\eta$ from noisy observations. To to so, it is useful to consider the TV denoising estimate $\hat{\eta}_{T}$. It is obtained by solving the optimization problem
@@
$\frac{1}{n}\|{\eta_{T}}-{{\hat{\eta}_{T}}}\|_{1}\leq\text{TVB}(\delta_{1}).$ (4)
@@
$\operatorname{CE}\leq\frac{1}{n}\sum_{i\in V}\left|s_{i}-\hat{\eta}(s_{i})\right|+\operatorname{BB}(|V|,\delta_{4},\hat{\sigma}^{2}_{\{|s_{i}-\hat{\eta}(s_{i})|\}_{i\in V}})+\operatorname{TVB}(\delta_{1})+\textnormal{PTB}(\delta_{2},\delta_{3}),$
@@ 5 Measuring Calibration Under Bounded Derivatives
The modification amounts to perturbing the probability output of the classifier by a small amount, specified by the perturbation bandwidth. This can be done either only at inference time or at both inference time and training time, to ensure better performance. We experimentally evaluate the effect of the perturbation in section 7 and conclude that there is almost no loss of performance relative to an unmodified classifier for realistic choices of the bandwidth. This allows us to write the calibration function of the perturbed classifier as
@@
with probability $1-\delta$ (picked so that $\delta_{1}+\delta_{2}=\delta$), with the randomness taken over the validation set, conditional on the training set, for a constant $R$ that depends [^5] For the truncated Epanechnikov kernel, we set $R=b_{1}h+\tfrac{1}{2}b_{2}h^{2}+\tfrac{1}{2}$ where $b_{1}$ and $b_{2}$ are uniform bounds on the first and second derivative of $\eta$. on the kernel $k$.
@@ 6 Related Work
Variants of bucketing were first introduced in meteorology (Murphy & Winkler, 1977) and have been used to approximate the calibration error ever since the ML community started paying attention to calibration (Zadrozny & Elkan, 2001; Niculescu-Mizil & Caruana, 2005; Naeini et al., 2015; Leathart et al., 2017; Guo et al., 2017; Vaicenavicius et al., 2019; Kumar et al., 2019). Nixon et al. (2019) has criticized vanilla bucketing and proposed alternative binning schemes. Maalej et al. (2025) used ECE-type binning in an empirical study of various calibration techniques.
@@
Dieye et al. (2025) have conducted an empirical study of several calibration metrics and concluded that the log-loss (or the Brier score) are better proxies for measuring calibration than a bucketing-based ECE estimate. The methods evaluated by Dieye et al. (2025) were all heuristics (they didn’t have non-asymptotic guarantees). Our work is different in that we provide certified upper bounds.
@@ 7 Experiments
The best way to evaluate an estimator is by comparing to the ground truth, which is unavailable for real-world datasets. We therefore introduced four synthetic examples of the calibration functions (see heading of Figure 3 for definitions). The figure shows the sample efficiency of bounding calibration error, defined as the gap between the probabilistic upper bound and the ground truth of the calibration error. Since we used synthetic datasets for this experiment, the ground truth is known and we can evaluate the gap exactly. We evaluate four principled methods: our kernel estimator (denoted as NW in the plot), our TV denoising estimator (denoted as TV) and Lipschitz bucketing (denoted as Lip+Bkt in the plot) as well as the ECE heuristic. We observe that all three principled methods are consistent (i.e. the error decreases with increasing dataset size), with NW having the best performance. Interestingly, the ECE heuristic is extremely competitive for the first three choices synthetic experiments, but fails completely for the fourth, in the sense that the error stays large and does not even decrease with increasing dataset size. This behavior is typical of heuristics – they sometimes work well but cannot be counted on for reliable estimation. We believe this underscores the importance of certified bounds.
@@
We repeated the all experiments on these plots 64 times and report mean performance. The figure 3 does not contain confidence bars because they are too small to be seen. This makes our results statistically significant.
@@ 11 Limitations
We provided full proofs for all our claims. We commit to release the source code that fully reproduces our figures upon acceptance.
@@ Appendix A Alternative Measures of Calibration Error
The $L_{q}$ calibration error [^8] We use the notation $L_{q}$ as opposed to the more common $L_{p}$ because we denoted a probability distribution with the symbol pp. provides a generalization of the $L_{1}$ calibration error by measuring deviations between predicted scores and true conditional probabilities using an $L_{q}$ norm. Formally, it is defined as $\mathbb{E}_{s}[|\eta(s)-s|^{q}]^{1/q}$, where $\eta(s)$ denotes the calibration function. By varying $q$, one can adjust the sensitivity to large deviations: higher values of $q$ emphasize larger miscalibration errors, while lower values treat all deviations more uniformly. Among these, the $\mathrm{L}_{1}$ version remains the most interpretable, as it directly quantifies the average absolute discrepancy between predicted probabilities and true event frequencies.
@@ B.2 Auxiliary Lemmas
$\sum_{k=1}^{j}\left(\frac{j-|T|}{|T|}\right)^{2}+\sum_{k=j+1}^{|T|}\left(\frac{j-|T|}{|T|}\right)^{2}\leq\frac{|T|}{4},$
@@ D.1.1 Hyperbolic secant kernel
Since $k(\cdot\mid s_{\text{orig}})\geq 0$ on $[0,1]$ and integrates to 11, it is a valid probability density function on $[0,1]$.
@@ Appendix E Implementation Details
Plug-in bandwidth
For the kernel $k^{\prime}$ we pick the bandwidth by minimizing a one–dimensional bias–variance surrogate obtained from the hh–dependent parts of the bound. With the Epanechnikov kernel, one can derive
$a=\tfrac{3}{8}b_{1},\ \ b=\tfrac{1}{10}b_{2},\ \ c=\tfrac{1.15}{2\sqrt{2n_{\text{tr}}}},$
so that optimizing $2b\,t^{5}+a\,t^{3}-\tfrac{1}{2}c=0$, gives the optimal $h=t^{2}$. We solve by a few Newton steps and clip to $10^{-4},1/4]$. This choice is an implementation detail; the theory only requires convex weights summing to one.
@@
We evaluate the Lipschitz+bucketing bound for every candidate partition defined by a bucket count $B$ and a fractional shift $r\in\{0,\dots,R\!-\!1\}$ (offset $r/B$); using a per-candidate confidence level $\delta/(BR)$, we take the minimum bound over all $(B,r)$. By a union bound, this aggregate is valid at overall level $\delta$.
@@ Appendix G Datasets
  • Spam Detection: A text dataset of emails labeled as either ”spam” or ”not-spam” Schulte et al. (2024). As for the IMDb data set, we fined-tuned a pre-trained BERT back-bone with a sequence classification head, using $1,000$ training examples, $250$ validation examples and $500$ test examples.
  • CIFAR-10: An image data set containing $10$ distinct classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship and truck (Krizhevsky, 2009). We make a binary task on this dataset by grouping means of transport (airplane, automobile, ship, truck) and animals (bird, cat, deer, dog, frog, horse). We fine-tune a pre-trained visual transformer (ViT) with a classification head. We use $2,500$ training examples, $250$ validation examples and $500$ test examples.
@@
  • Phishing: A dataset of websites’ URLs with binary phishing and not-phishing labels (aisuko, 2024). We fine-tuned a BERT-based classifier model with $10,000$ training examples and $250$ validation examples. We then use $500,000$ examples as test.
  • Civil Comments: A large data set of text comments from an archive of the Civil Comments platform; a commenting plugin for independent news sites (Borkan et al., 2019). In the original set, the comments have a series of scores assigned to them, delving into different aspects of harmfulness and toxicity. We make a binary data set based on the parent toxicity score; if this score is greater than 0, the label is 11 (toxicity present) and 0 otherwise. As for the phishing data set, we use $10,000$ training examples and $250$ validation examples to fine-tune a BERT-based text classifier. We use on 3.83.8M as test set.
  • Yelp Polarity: A data set of Yelp reviews, where the star ratings have been binarised ($\geq 3$ stars is considered as positive) (Zhang et al., 2015). We fine-tuned a BERT-based classifier model with $10,000$ training examples and $250$ validation examples. We then use $560,000$ examples as test.

### v3

@@ Abstract
We make two contributions to the problem of estimating the $L_{1}$ calibration error of a binary classifier from a finite dataset. First, we provide an upper bound on the calibration error for any classifier where the calibration function has bounded variation. Second, we provide a method of modifying any classifier so that its calibration error can be upper bounded efficiently without significantly impacting classifier performance and without any restrictive assumptions. All our results are non-asymptotic and distribution-free. We conclude by providing advice on how to measure calibration error in practice. Our methods yield practical procedures that can be run on real-world datasets with modest overhead.
@@ 3 Structure of Results
Here, we used the symbol BB to denote the concentration term in the Bernstein bound. We could have equally used Hoeffding inequality here, but Bernstein is slightly sharper.
@@ 4 Measuring Calibration Under Bounded Variation
To formalize our assumption on $\eta$, we will need the concept of a Bounded Variation space (DeVore & Lorentz, 1993). For a function $f$, we can define total variation on the interval $I$ as
@@
where the supremum is taken over the set of all finite sequences $s_{1},s_{2},\dots,s_{n}$ contained in the interval $I$. We say that the function $f$ is of bounded variation, written $f\in\text{BV}([0,1])$ when $\text{TV}(f,[0,1])<\infty$. Our theoretical work in this section is based on the assumption that the calibration function $\eta$ satisfies $\text{TV}(\eta,[0,1])\leq V$ for some known constant $V$.
@@
Under the bounded variation assumption, we want to reconstruct $\eta$ from noisy observations. To do so, it is useful to consider the TV denoising estimate $\hat{\eta}_{T}$. It is obtained by solving the optimization problem
@@
$\frac{1}{|T|}\|{\eta_{T}}-{{\hat{\eta}_{T}}}\|_{1}\leq\text{TVB}(\delta_{1}).$ (4)
@@
$\operatorname{CE}\leq\frac{1}{|T|}\sum_{i\in V}\left|s_{i}-\hat{\eta}(s_{i})\right|+\operatorname{BB}(|V|,\delta_{4},\hat{\sigma}^{2}_{\{|s_{i}-\hat{\eta}(s_{i})|\}_{i\in V}})+\operatorname{TVB}(\delta_{1})+\textnormal{PTB}(\delta_{2},\delta_{3}),$
@@ 5 Measuring Calibration Under Bounded Derivatives
The modification amounts to perturbing the probability output of the classifier by a small amount, specified by the perturbation bandwidth. This can be done either only at inference time or at both inference time and training time, to ensure better performance. We experimentally evaluate the effect of the perturbation in section 7 and conclude that there is almost no loss of performance relative to an unmodified classifier for realistic choices of the bandwidth. The calibration function of the perturbed classifier can be written as
@@
with probability $1-\delta$ (picked so that $\delta_{1}+\delta_{2}=\delta$), with the randomness taken over the validation set, conditional on the training set, for a constant $R$ that depends [^5] For the truncated Epanechnikov kernel, we set $R=b_{1}h+\tfrac{1}{2}b_{2}h^{2}+\tfrac{1}{2}$ where $b_{1}$ and $b_{2}$ are uniform bounds on the first and second derivative of $\eta$. on the kernel $k^{\prime}$.
@@ 6 Related Work
Variants of bucketing were first introduced in meteorology (Murphy & Winkler, 1977) and have been used to approximate the calibration error ever since the ML community started paying attention to calibration (Zadrozny & Elkan, 2001; Niculescu-Mizil & Caruana, 2005; Naeini et al., 2015; Leathart et al., 2017; Guo et al., 2017; Vaicenavicius et al., 2019; Kumar et al., 2019). Nixon et al. (2019) has criticized vanilla bucketing and proposed alternative binning schemes.
@@
Dieye et al. (2025) have conducted an empirical study of several calibration metrics and concluded that the log-loss (or the Brier score) are better proxies for measuring calibration than a bucketing-based ECE estimate. The methods evaluated by Dieye et al. (2025) were all heuristics (they didn’t have non-asymptotic guarantees). Maalej et al. (2025) used ECE-type binning in an empirical study of various calibration techniques. Our work is different in that we provide certified upper bounds.
@@ 7 Experiments
The best way to evaluate an estimator is by comparing to the ground truth, which is unavailable for real-world datasets. We therefore introduced four synthetic examples of the calibration functions (see heading of figure 3 for definitions). The figure shows the sample efficiency of bounding calibration error, defined as the gap between the probabilistic upper bound and the ground truth of the calibration error. Since we used synthetic datasets for this experiment, the ground truth is known and we can evaluate the gap exactly. We evaluate four principled methods: our kernel estimator (denoted as NW in the plot), our TV denoising estimator (denoted as TV) and Lipschitz bucketing (denoted as Lip+Bkt in the plot) as well as the ECE heuristic. We observe that all three principled methods are consistent (i.e. the error decreases with increasing dataset size), with NW having the best performance. Interestingly, the ECE heuristic is extremely competitive for the first three choices of synthetic experiments, but fails completely for the fourth, in the sense that the error stays large and does not even decrease with increasing dataset size. This behavior is typical of heuristics – they sometimes work well but cannot be counted on for reliable estimation. We believe this underscores the importance of certified bounds.
@@
We repeated the all experiments on these plots 64 times and report mean performance. Figure 3 does not contain confidence bars because they are too small to be seen. This makes our results statistically significant.
@@ 11 Limitations
We provided full proofs for all our claims. Source code is available at https://github.com/spotify-research/calibration.
@@ Appendix A Alternative Measures of Calibration Error
The $L_{q}$ calibration error [^8] We use the notation $L_{q}$ as opposed to the more common $L_{p}$ because we denoted a probability distribution with the symbol $p$. provides a generalization of the $L_{1}$ calibration error by measuring deviations between predicted scores and true conditional probabilities using an $L_{q}$ norm. Formally, it is defined as $\mathbb{E}_{s}[|\eta(s)-s|^{q}]^{1/q}$, where $\eta(s)$ denotes the calibration function. By varying $q$, one can adjust the sensitivity to large deviations: higher values of $q$ emphasize larger miscalibration errors, while lower values treat all deviations more uniformly. Among these, the $L_{1}$ version remains the most interpretable, as it directly quantifies the average absolute discrepancy between predicted probabilities and true event frequencies.
@@ B.2 Auxiliary Lemmas
$\sum_{k=1}^{j}\left(\frac{j-|T|}{|T|}\right)^{2}+\sum_{k=j+1}^{|T|}\left(\frac{j}{|T|}\right)^{2}\leq\frac{|T|}{4},$
@@ B.5 Bounded Variation Population Transfer
Note that the term $\sqrt{\frac{1}{2n}\log\frac{2}{\delta}}$ in lemma 5 is only needed if $F$ assigns nonzero probability mass to the zero score.
@@ D.1.1 Hyperbolic secant kernel
Since $k(\cdot\mid s_{\text{orig}})\geq 0$ on $[0,1]$ and integrates to $1$, it is a valid probability density function on $[0,1]$.
@@ Appendix E Implementation Details
Plug-in bandwidth.
We pick $h^{\prime}$ (the bandwidth of $k^{\prime}$) by minimizing a one-dimensional proxy built from the $h^{\prime}$-dependent terms in the smoothing-error bound (Lemma 6): the bias terms are $b_{1}\sum_{i\in T}w_{i}(s^{\prime})\,|s^{\prime}-s_{i}|$ and $\tfrac{b_{2}}{2}\sum_{i\in T}w_{i}(s^{\prime})\,(s^{\prime}-s_{i})^{2}$, and the stochastic term is $\tfrac{1}{2}\sqrt{\sum_{i\in T}w_{i}(s^{\prime})^{2}}$. Approximating these sums by their continuous Epanechnikov-kernel moments for locally uniform score density (ignoring boundary effects), with $K(u)=\tfrac{3}{4}(1-u^{2})\mathbf{1}\{|u|\leq 1\}$ and $u=(s^{\prime}-s_{i})/h^{\prime}$, yields $\sum_{i}w_{i}|s^{\prime}-s_{i}|\approx h^{\prime}\!\int|u|K(u)\,du=\tfrac{3}{8}h^{\prime}$ and $\sum_{i}w_{i}(s^{\prime}-s_{i})^{2}\approx h^{\prime 2}\!\int u^{2}K(u)\,du=\tfrac{1}{5}h^{\prime 2}$, hence $a=\tfrac{3}{8}b_{1}$ and $b=\tfrac{1}{10}b_{2}$ in the proxy
$f(h^{\prime})=a\,h^{\prime}+b\,{h^{\prime}}^{2}+\frac{c}{\sqrt{h^{\prime}}}.$
For the stochastic term we use the standard approximation $\sqrt{\sum_{i}w_{i}^{2}}\approx\sqrt{\|K\|_{2}^{2}/(|T|h^{\prime})}$ (with $\|K\|_{2}^{2}=\int K(u)^{2}\,du=3/5$ for Epanechnikov) and take a mildly conservative constant in implementation, giving $c=\tfrac{1.15}{2\sqrt{2|T|}}$. Setting $f^{\prime}(h^{\prime})=0$ gives $a+2bh^{\prime}-\tfrac{1}{2}c\,{h^{\prime}}^{-3/2}=0$; multiplying by ${h^{\prime}}^{3/2}$ and substituting $t=\sqrt{h^{\prime}}$ yields the quintic $2b\,t^{5}+a\,t^{3}-\tfrac{1}{2}c=0$, and we recover the bandwidth as $h^{\prime}=t^{2}$. We solve by a few Newton steps and clip to $10^{-4},1/4]$. This choice is an implementation detail; the theory only requires convex weights summing to one.
@@
Here, “bucketing” means uniform bins in score space: for a given bucket count $B$ we use an (approximately) equal-width grid over $[0,1]$ with bin width $\approx 1/B$, and the “shift” moves the internal bin boundaries by a small fraction of one bin width (while keeping the endpoints at 0 and $1$). This is not quantile/equal-count binning. We then evaluate the Lipschitz+bucketing bound for every candidate partition defined by a bucket count $B$ and a fractional shift $r\in\{0,\dots,R\!-\!1\}$ (offset $r/B$); using a per-candidate confidence level $\delta/(BR)$, we take the minimum bound over all $(B,r)$. By a union bound, this aggregate is valid at overall level $\delta$.
@@ Appendix G Datasets
  • Spam Detection: A text dataset of emails labeled as either ‘spam’ or ‘not-spam’ Schulte et al. (2024). As for the IMDb data set, we fined-tuned a pre-trained BERT back-bone with a sequence classification head, using $1,000$ training examples, $250$ validation examples and $500$ test examples.
  • CIFAR-10: An image data set containing $10$ distinct classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship and truck (Krizhevsky, 2009). We make a binary task on this dataset by grouping means of transport (airplane, automobile, ship, truck) and animals (bird, cat, deer, dog, frog, horse). We fine-tune a pre-trained visual transformer (ViT) with a classification head. We use $2500$ training examples, $250$ validation examples and $500$ test examples.
@@
  • Phishing: A dataset of websites’ URLs with binary phishing and not-phishing labels (Li, 2024). We fine-tuned a BERT-based classifier model with $10000$ training examples and $250$ validation examples. We then use $500000$ examples as test.
  • Civil Comments: A large data set of text comments from an archive of the Civil Comments platform; a commenting plugin for independent news sites (Borkan et al., 2019). In the original set, the comments have a series of scores assigned to them, delving into different aspects of harmfulness and toxicity. We make a binary data set based on the parent toxicity score; if this score is greater than 0, the label is $1$ (toxicity present) and 0 otherwise. As for the phishing data set, we use $10,000$ training examples and $250$ validation examples to fine-tune a BERT-based text classifier. We use on 3.83.8M as test set.
  • Yelp Polarity: A data set of Yelp reviews, where the star ratings have been binarised ($\geq 3$ stars is considered as positive) (Zhang et al., 2015). We fine-tuned a BERT-based classifier model with $10000$ training examples and $250$ validation examples. We then use $560000$ examples as test.